# Integral canonical models of exceptional Shimura varieties

### v1

@@ 8.1 The Fontaine-Laffaille module on complete local rings
Consider the subgroups of $\operatorname{GL}(_{{}_{\mathrm{cris}}}\mathbb{V}_{\tilde{x}})$ and $\operatorname{GL}(_{{}_{\mathrm{cris}}}\mathbb{V}^{x})$ that fix the tensors $s_{\alpha}$ and the filtrations ${\mathrm{Fil}}^{\bullet}$. The existence of the integral model of $\mathscr{P}/\mathscr{S}_{K}(G,X)$ implies that these are parabolic subgroups $P_{W}\subset G_{W}$ and $P_{R^{x}}\subset G_{R^{x}}$ respectively – the identification $\iota_{\tilde{x}}$ and $\iota^{x}$ in Corollaries 8.3 and 8.4 allow us to think of the parabolic subgroups as subgroups of $G_{W}$ and $G_{R^{x}}$. Further, $P_{R^{x}}$ clearly specializes to $P_{W}$ via the point $\tilde{x}$. Consider now the two parabolic subgroups $P_{R^{x}}$ and $P_{W}\otimes R^{x}$ of $F_{R^{x}}$. These two subgroups are equal (and therefore conjugate) at the unique closed point. By [Conrad:ReductiveSchemes, Corollary 5.2.7], the two parabolics are $\mathcal{G}(R^{x})$-conjugate. Conjugating $\iota$ by this element still respects the tensors (which are fixed by elements of $G$) and now respects the filtrations by construction. The proposition follows. ∎
@@ 8.1.3 The Fontaine-Laffaille module on complete local rings
We will now prove the main result of this section, which is to give a description of the Fontaine-Laffaille data on $S_{K}(G,X)^{x}$ analogous to the case of abelian type Shimura varieties following Kisin [Kisinintegral]. Let notation be as in the previous subsections. Recall that we have chosen an isomorphism ${}_{{}_{\mathrm{cris}}}\mathbb{V}_{\tilde{x}}\otimes R^{x}\rightarrow_{{}_{\mathrm{cris}}}\mathbb{V}^{x}$ of filered modules compatible with the $G$-structure defined in Subsection 8.1.1. We note that such an isomorphism is unique only up to an element of $P(R^{x})$, where $P$ is the parabolic subgroup of $G$ defined by the filtration. Also recall that we have chosen a $G$-valued co-character (defined over $W$) that splits ${\mathrm{Fil}}^{\bullet}$. Let $U^{\operatorname{opp}}$ denote the opposite unipotent of $\mu$, and let $\widehat{U}^{\operatorname{opp}}$ denote the completion of $U^{\operatorname{opp}}$ at the identity mod $p$.
@@ 9 The $\mu$-ordinary locus: CM lifts and Tate’s isogeny theorem
We keep the notation of the previous section. Let $B\subset G$ be a Borel subgroup defined over $\mathbb{Z}_{p}$, and let $T\subset B\subset G$ be a maximal torus that splits over an etale extension of $\mathbb{Z}_{p}$(that $G/\mathbb{Z}_{p}$ is reductive implies that these objects exist). Let $x\in\mathscr{S}_{K}(G,X)(\mathbb{F})$. Associated to the $F$-crystal ${}_{{}_{\mathrm{cris}}}\mathbb{V}_{x}$ is the Frobenius element $b_{x}\in G(L)$, which is well defined up to $\sigma$-conjugation by $G(W)$. Let $\nu_{x}\in X_{*}(T)$ be the dominant element that is conjugate to $\nu_{b_{x}}$, the Newton cocharacter associated to $b_{x}$ (see work of Kottwitz [Kott1] and [Kott2] for definitions and more details). Note that the conjugacy class of $\nu_{b_{x}}$ depends only on the $\sigma$-conjugacy class of $b_{x}$, and so the definition of $\nu_{x}$ is independent of the choice of representative of the $\sigma$-conjugacy class of $b_{x}$. We note that $\nu_{x}$ is defined over $\mathbb{Z}_{p}$. Let $\mu_{x}\in X_{*}(T)$ denote the dominant co-character with $b_{x}\in G(W)\sigma(\mu_{x})(p)G(W)$. The co-character $\mu_{x}$ might not be defined over $\mathbb{Z}_{p}$, and let $\bar{\mu}_{x}=\frac{1}{[E_{p}:\mathbb{Q}_{p}]}\sum_{\tau\in\operatorname{Gal}(E_{p}/\mathbb{Q}_{p})}\tau(\sigma)$ denote the Galois average of $\mu$, where $E_{p}/\mathbb{Q}_{p}$ is some finite unramified extension splitting $T$. We have that $\nu_{x}\preceq\bar{\mu_{x}}$ (for example, see [Gashi]), where two dominant fractional cocharacters satisfy $\mu^{\prime}\preceq\mu^{\prime\prime}$ if and only if $\mu^{\prime\prime}-\mu^{\prime}$ can be written as a non-negative rational linear combination of positive co-roots. We have the following definition.
@@
There exists $g\in G(W)$ such that $g^{-1}b\sigma(g)=(\sigma\mu)(p)$.
We now change coordinates on (by $\sigma$-conjugating by an element of $G(W)$) and we may assume that $b=\sigma(\mu)p$. The following proposition defines the canonical lift of a $\mu$-ordinary point.
@@ 9.1 Non-emptiness of the $\mu$-ordinary locus
By Propositions 9.10 and 9.12, it suffices to find a Weyl special point $(T,h)$ where $T$ is quasi-split at $p$. But this follows directly from [ChaiOortJacobians, Proposition 5.11] by choosing $\Sigma_{1}=\{p,\infty\}$, $U_{\infty}$ as in Remark 5.12 of loc. cit., and $U_{p}$ to be the set of set of elements in $\operatorname{Lie}G(\mathbb{Z}_{p})$ which reduce to an element of $\operatorname{Lie}G(\mathbb{Z}_{p})$ whose stabilizer in $G_{\mathbb{F}_{p}}$ is equal to a maximal torus containing a maximal split torus. ∎

### v2

@@ 7.1 The Fontaine-Laffaille module on complete local rings
Consider the subgroups of $\operatorname{GL}(_{{}_{\mathrm{cris}}}\mathbb{V}_{\tilde{x}})$ and $\operatorname{GL}(_{{}_{\mathrm{cris}}}\mathbb{V}^{x})$ that fix the tensors $s_{\alpha}$ and the filtrations ${\mathrm{Fil}}^{\bullet}$. The existence of the integral model of $\mathscr{P}/\mathscr{S}_{K}(G,X)$ implies that these are parabolic subgroups $P_{W}\subset G_{W}$ and $P_{R^{x}}\subset G_{R^{x}}$ respectively – the identification $\iota_{\tilde{x}}$ and $\iota^{x}$ in Corollaries 7.3 and 7.4 allow us to think of the parabolic subgroups as subgroups of $G_{W}$ and $G_{R^{x}}$. Further, $P_{R^{x}}$ clearly specializes to $P_{W}$ via the point $\tilde{x}$. Consider now the two parabolic subgroups $P_{R^{x}}$ and $P_{W}\otimes R^{x}$ of $F_{R^{x}}$. These two subgroups are equal (and therefore conjugate) at the unique closed point. By [Conrad:ReductiveSchemes, Corollary 5.2.7], the two parabolics are $\mathcal{G}(R^{x})$-conjugate. Conjugating $\iota$ by this element still respects the tensors (which are fixed by elements of $G$) and now respects the filtrations by construction. The proposition follows. ∎
@@ 7.1.3 The Fontaine-Laffaille module on complete local rings
We will now prove the main result of this section, which is to give a description of the Fontaine-Laffaille data on $S_{K}(G,X)^{x}$ analogous to the case of abelian type Shimura varieties following Kisin [Kisinintegral]. Let notation be as in the previous subsections. Recall that we have chosen an isomorphism ${}_{{}_{\mathrm{cris}}}\mathbb{V}_{\tilde{x}}\otimes R^{x}\rightarrow_{{}_{\mathrm{cris}}}\mathbb{V}^{x}$ of filtered modules compatible with the $G$-structure defined in Subsection 7.1.1. We note that such an isomorphism is unique only up to an element of $P(R^{x})$, where $P$ is the parabolic subgroup of $G$ defined by the filtration. Also recall that we have chosen a $G$-valued co-character (defined over $W$) that splits ${\mathrm{Fil}}^{\bullet}$. Let $U^{\operatorname{opp}}$ denote the opposite unipotent of $\mu$, and let $\widehat{U}^{\operatorname{opp}}$ denote the completion of $U^{\operatorname{opp}}$ at the identity mod $p$.
@@ 8 The $\mu$-ordinary locus: CM lifts and Tate’s isogeny theorem
We keep the notation of the previous section. Let $B\subset G$ be a Borel subgroup defined over $\mathbb{Z}_{p}$, and let $T\subset B\subset G$ be a maximal torus that splits over an etale extension of $\mathbb{Z}_{p}$(that $G/\mathbb{Z}_{p}$ is reductive implies that these objects exist). Let $x\in\mathscr{S}_{K}(G,X)(\mathbb{F})$. Associated to the $F$-crystal ${}_{{}_{\mathrm{cris}}}\mathbb{V}_{x}$ is the Frobenius element $b_{x}\in G(L)$, which is well defined up to $\sigma$-conjugation by $G(W)$. Let $\nu_{x}\in X_{*}(T)$ be the dominant element that is conjugate to $\nu_{b_{x}}$, the Newton cocharacter associated to $b_{x}$ (see work of Kottwitz [Kott1] and [Kott2] for definitions and more details). Note that the conjugacy class of $\nu_{b_{x}}$ depends only on the $\sigma$-conjugacy class of $b_{x}$, and so the definition of $\nu_{x}$ is independent of the choice of representative of the $\sigma$-conjugacy class of $b_{x}$. We note that $\nu_{x}$ is defined over $\mathbb{Z}_{p}$. Let $\mu_{x}\in X_{*}(T)$ denote the dominant co-character with $b_{x}\in G(W)\sigma(\mu_{x})(p)G(W)$. The co-character $\mu_{x}$ might not be defined over $\mathbb{Z}_{p}$, and let $\bar{\mu}_{x}=\frac{1}{[E_{p}:\mathbb{Q}_{p}]}\sum_{\tau\in\operatorname{Gal}(E_{p}/\mathbb{Q}_{p})}\tau(\mu_{x})$ denote the Galois average of $\mu$, where $E_{p}/\mathbb{Q}_{p}$ is some finite unramified extension splitting $T$. We have that $\nu_{x}\preceq\bar{\mu_{x}}$ (for example, see [Gashi]), where two dominant fractional cocharacters satisfy $\mu^{\prime}\preceq\mu^{\prime\prime}$ if and only if $\mu^{\prime\prime}-\mu^{\prime}$ can be written as a non-negative rational linear combination of positive co-roots. We have the following definition.
@@
Let $x\in\mathscr{S}_{K}(G,X)(\mathbb{F})$ be a $\mu$-ordinary point with $b_{x},\nu_{x}$ and $\mu_{x}$ as above. Then there exists $g\in G(W)$ such that $g^{-1}b_{x}\sigma(g)=(\sigma\mu_{x})(p)$.
We now change coordinates on (by $\sigma$-conjugating by an element of $G(W)$) and we may assume that $b_{x}=\sigma(\mu_{x})p$. The following proposition defines the canonical lift of a $\mu$-ordinary point.
@@ 8.1 Non-emptiness of the $\mu$-ordinary locus
By Propositions 8.10 and 8.12, it suffices to find a Weyl special point $(T,h)$ where $T$ is quasi-split at $p$. But this follows directly from [ChaiOortJacobians, Proposition 5.11] by choosing $\Sigma_{1}=\{p,\infty\}$, $U_{\infty}$ as in Remark 5.12 of loc. cit., and $U_{p}$ to be the set of set of elements in $\operatorname{Lie}G(\mathbb{Z}_{p})$ which reduce to an element of $\operatorname{Lie}G(\mathbb{Z}_{p})$ whose stabilizer in $G_{\mathbb{F}_{p}}$ is equal to a maximal torus containing a maximal split torus. ∎